# Infrequent Resolving Algorithm for Online Linear Programming

### v3

@@ 2.4 Finite Resolving
$\mathcal{T}^{F}_{A}(M)=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-2}}\right\rceil\right\}.$
To facilitate understanding, we illustrate the resolving schedule in Figure 2. As Figure 2 shows, when the number of resolvings is restricted, we should invest more computational power to the approximation set, and the learning set only needs two time points. More specifically, we will not solve the fluid model until we get enough samples to derive a relatively accurate estimation of $\boldsymbol{p}$, i.e., at the first time point $\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil$. Then, at time point $\lceil T/2\rceil$, we solve the fluid model with a more accurate estimation based on the collected samples. As for the approximation set, in order to correct the approximation error in time, the smaller time points (further away from $T$) are more crucial. Therefore, we keep smaller time points when the number of resolvings is limited.
@@
For $M\geq 2$, given the finite-resolving schedule $\mathcal{T}^{F}(M)$ with $\beta\in(\frac{1}{2},1)$ the regret of the AIR policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}})$.
@@ 3 Known Arrival Probabilities
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
Given the arrival probabilities and the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
@@ 3.1 Finite Resolving
According to Theorem 4, by setting $\beta=1/2+\epsilon$, we can guarantee an $\mathcal{O}(T^{(1/2+\epsilon)^{M}})\approx\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret by solving LPs $M$ times. The result is close to Theorem 6.1 in Jasin and Kumar (2012), which shows that $M$ LP solving times can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound under the non-degeneracy assumption. However, different from Jasin and Kumar (2012), our resolving schedule also works for the degenerate case.
@@ 4.1 OLP Policy Comparison
In the following, we compare the performance of the above algorithms. For each parameter set, we run 2,000 simulations. To gain deeper insights, we first focus on the single resource problem similar to the one studied in Bumpensanti and Wang (2020). Specifically, we consider two types of customers and one type of resource. The arrival probabilities of either customer type is $0.5$. The rewards of the two types of customers are $r_{1}=2$ and $r_{2}=1$, and each customer consumes one unit of resource. The budget per period is $\rho$.
First, in Figure 4, we fix the time horizon $T$ as $50,000$, and test the algorithms as the budget factor $\rho$ changes. According to Figure 4, the regret of most policies is large when $\rho$ is close to 0.5. Due to the setup of the single-resource problem, the fluid problem at time $1$ is degenerate when $\rho=0.5$. Therefore, the degenerate case is the hardest case for most policies. However, both the AIR policy and the AFR policy guarantee a constant regret in this case.
@@ B.7 Proof of Theorem 2
    For the time interval $\left[1,\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$, Algorithm 1 incurs at most $T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}$ regret.
@@ B.12 Proof of Theorem 4
B.13 Proof of Lemma 3

### v4

@@ 2.4 Finite Resolving
$\mathcal{T}^{F}_{A}(M)=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-2}}\right\rceil\right\},$
where $\epsilon$ can be any positive constant. To facilitate understanding, we illustrate the resolving schedule in Figure 2. As Figure 2 shows, when the number of resolvings is restricted, we should invest more computational power to the approximation set, and the learning set only needs two time points. More specifically, we will not solve the fluid model until we get enough samples to derive a relatively accurate estimation of $\boldsymbol{p}$, i.e., at the first time point $\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil$. Then, at time point $\lceil T/2\rceil$, we solve the fluid model with a more accurate estimation based on the collected samples. The insight behind this choice is that, at early stages, there is plenty of inventory and wrong actions can be made up by the remaining periods. Specifically, if we accept too much type-$j$ customers at early periods, then we can reject more in the remaining periods to make up. However, at late periods, we cannot tolerate such many wrong actions. As for the approximation set, in order to correct the approximation error in time, the smaller time points (further away from $T$) are more crucial. Therefore, we keep smaller time points when the number of resolvings is limited.
@@
For $M\geq 2$ and any $\epsilon>0$, given the finite-resolving schedule $\mathcal{T}^{F}(M)$ with $\beta\in(\frac{1}{2},1)$ the regret of the AIR policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}})$.
@@
Remark 2 (Role of Finite-Support Assumption).
As mentioned in the introduction, we assume that the underlying distribution $\mathcal{P}$ has finite supports. We would like to explain the role of the assumption in the above proof. First, the finite-support assumption enables us to aggregate customers by types and derive a tractable fluid model (4) with finite variables. If the distribution $\mathcal{P}$ has infinite supports, then the fluid model becomes an intractable infinite linear program. In that case, the typical approach is to work on the dual problem instead of the primal (see, Li and Ye 2022, Bray 2024 and Ma et al. 2024). However, in the degenerate case, the linkage between the optimal primal solution and the optimal dual solution becomes relatively weak, resulting in large regrets of the above algorithms.
Second, the finite-support assumption ensures that the arrival probability of each customer type is positive, which plays an important role in Proposition 4. As we explained after Proposition 4, one of the key parts is to deduce $\frac{T-t+1}{4}p_{j}\geq 1$ when $T-t+1$ is greater than a constant independent of $T$. However, for the continuous-support case, the probability measure of each type is zero and hence the proof cannot hold.
Indeed, as Bray (2024) showed, under the continuous-support assumption, the regret of any policy is lower bounded by $\Omega(\log T)$, i.e., a constant-regret algorithm is impossible. ∎
@@ 3 Known Arrival Probabilities
     if  $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$, $u_{j}^{t}>1$, and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
For any $\epsilon>0$, given the arrival probabilities and the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
@@ 3.1 Finite Resolving
According to Theorem 4, by setting $\beta=1/2+\epsilon$, we can guarantee an $\mathcal{O}(T^{(1/2+\epsilon)^{M}})\approx\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret by solving LPs $M$ times. The result is close to Theorem 6.1 in Jasin and Kumar (2012), which shows that $M$ LP solving times can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound under the non-degeneracy assumption. However, different from Jasin and Kumar (2012), our resolving schedule also works for the degenerate case. Moreover, in the following proposition, we provide a lower bound for the regret of the AIR-KP policy with any two resolving time points.
Proposition 5.
Given any resolving schedule with no more than two resolvings, the regret of the AIR-KP policy is $\Omega(T^{1/4})$.
According to Theorem 4 and Proposition 5, the regret bound $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ (with $\epsilon\to 0$) is nearly tight when $M=2$.
@@ 4.1 OLP Policy Comparison
In the following, we compare the performance of the above algorithms. [^4] Note that the distribution assumptions and the non-degeneracy assumption in the Gao et al. (2024) and Ma et al. (2024) may not be satisfied in the numerical experiments. For each parameter set, we run 2,000 simulations. To gain deeper insights, we first focus on the single resource problem similar to the one studied in Bumpensanti and Wang (2020). Specifically, we consider two types of customers and one type of resource. The arrival probabilities of either customer type is $0.5$. The rewards of the two types of customers are $r_{1}=2$ and $r_{2}=1$, and each customer consumes one unit of resource. The budget per period is $\rho$.
First, in Figure 4, we fix the time horizon $T$ as $50,000$, and test the algorithms as the budget factor $\rho$ changes. Due to the setup of the single-resource problem, the fluid problem at time $1$ is degenerate when $\rho$ takes the value $0.5$ and hence the non-degeneracy measure $\delta$ equals $|\rho-0.5|$. According to Figure 4, the regret of most policies is large when $\rho$ is close to 0.5 (i.e, $\delta\to 0$). Therefore, the degenerate case is the hardest case for most policies. However, both the AIR policy and the AFR policy guarantee a constant regret in this case.
@@
Then, we investigate the impact of the hyper-parameters (i.e., $\alpha$ and $\beta$) on the regret in Table 5. According to Table 5, as $\alpha$ or $\beta$ increases, the regret first decreases steeply and then gradually becomes steady. Therefore, to implement our policy in practice, it is recommended to set slightly large hyper-parameters ($\alpha$ and $\beta$).
$\alpha$ Regret # of Resolvings
0.15 887.4 10
0.25 211.7 10
0.35 79.6 11
0.45 13.6 11
0.55 2.4 12
0.65 2.2 14
0.75 2.0 16
0.85 2.0 22
0.95 1.8 45
(a) Fix $\beta=0.7$
$\beta$ Regret # of Resolvings
0.55 9.2 12
0.60 5.6 13
0.65 3.3 14
0.70 2.1 15
0.75 1.9 16
0.80 1.8 18
0.85 1.8 22
0.90 1.7 28
0.95 1.7 45
(b) Fix $\alpha=0.7$
Table 5: Regret of AIR Policy as $\alpha$ or $\beta$ Changes ($T=30,000$).
@@ B.7 Proof of Theorem 2
    For the time interval $\left[1,\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$, Algorithm 1 rejects all customers and incurs at most $T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}$ regret.
@@ B.12 Proof of Theorem 4
B.13 Proof of Proposition 5
According to Theorem 3, we have $\mathbb{E}\left[\phi(T\boldsymbol{\rho},\boldsymbol{Z}^{1})\right]-V^{*}(T)=O(1)$, i.e., the difference between the hindsight problem and the optimal problem is upper bounded by a constant. Thus, in the following, we will compare the policy with the hindsight problem.
Let $T_{1}$ and $T_{2}$ (with $T_{1}<T_{2}$) denote the two resolving time points. Similar to Bumpensanti and Wang (2020), in this proof, we consider two types of customers and one type of resource. The arrival probabilities of either customer type is $0.5$. The rewards of the two types of customers are $r_{1}=2$ and $r_{2}=1$, and each customer consumes one unit of resource. The budget per period is $0.5$, resulting in a degenerate case.
We start with the case where $T_{1}=\omega(T^{1/4})$, which means that $\lim_{T\to\infty}\frac{T_{1}}{T^{1/4}}=\infty$. In this case, due to the initialization in Algorithm 2, the algorithm will reject all customers during the time interval $[1,T_{1})$. Let $\Lambda_{j}(t_{1},t_{2})$ denote the random number of type-$j$ arrivals during the time interval $[t_{1},t_{2}]$. We consider the event where $\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}\leq\Lambda_{1}(1,T_{1}-1)\leq\frac{1}{2}(T_{1}-1)$ and $\Lambda_{1}(T_{1},T)\leq\frac{1}{2}(T-T_{1}+1)$. In this case, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{1}-1)+\Lambda_{1}(T_{1},T)\leq\frac{1}{2}T$ and hence the hindsight problem will accept all type-1 customers. However, the algorithm rejects $\Lambda_{1}(1,T_{1}-1)$ type-1 customers, and hence the corresponding revenue loss is at least $\Lambda_{1}(1,T_{1}-1)\geq\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}=\Omega(T^{1/4})$. Then, similar to Bumpensanti and Wang (2020), according to the Berry-Esseen theorem (see Shevtsova 2011), we bound the probability of the event as follows:
$\mathbb{P}\left(\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}\leq\Lambda_{1}(1,T_{1}-1)\leq\frac{1}{2}(T_{1}-1)\;\&\;\Lambda_{1}(T_{1},T)\leq\frac{1}{2}(T-T_{1}+1)\right)$
$=$ $\mathbb{P}\left(-\frac{\sqrt{T_{1}-1}}{\frac{1}{2}\sqrt{T_{1}-1}}\leq\frac{\Lambda_{1}(1,T_{1}-1)-\frac{1}{2}(T_{1}-1)}{\frac{1}{2}\sqrt{T_{1}-1}}\leq 0\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{1},T)-\frac{1}{2}(T-T_{1}+1)}{\frac{1}{2}\sqrt{T-T_{1}+1}}\leq 0\right)$
$\geq$ $\left(0.477-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T_{1}-1}}\right)\cdot\left(0.5-\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{1}+1}}\right),$
which is greater than a constant when $T$ is no less than a threshold. Therefore, the regret will be the order of $\Omega(T^{1/4})$. Then, we consider the case where $T_{1}=O(T^{1/4})$, and study different choices of $T_{2}$.
Case I: There exists a constant $c_{12}<1$ such that $T_{1}\leq T_{2}\leq c_{12}\cdot T$.
In this case, we consider the event when $\frac{1}{2}(T_{2}-1)\leq\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)+\sqrt{T_{2}-1}$ and $\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)-\sqrt{T}$. Under this event, the total number of type-1 arrivals is $\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}T+\sqrt{T_{2}-1}-\sqrt{T}\leq\frac{1}{2}T-(1-\sqrt{c_{12}})\sqrt{T}$, and hence the hindsight optimal policy accepts at least $(1-\sqrt{c_{12}})\sqrt{T}$ type-2 customers. Then, we analyze the performance of the AIR-KP policy with the resolving schedule $\{T_{1},T_{2}\}$. During the time interval $[1,T_{1})$, the policy rejects all customers. Then, the algorithm resolves the fluid model and derives the solution $\boldsymbol{u}^{T_{1}}=\left(\frac{1}{2}(T-T_{1}),\frac{1}{2}T_{1}\right)$ and the demand estimation $\boldsymbol{d}^{T_{1}}=\left(\frac{1}{2}(T-T_{1}),\frac{1}{2}(T-T_{1})\right)$. Thus, during the interval $[T_{1},T_{2})$, the algorithm will accept all type-1 customers and no greater than $\frac{1}{2}T_{1}$ type-2 customers, resulting in $\nu\geq\Lambda_{1}(1,T_{2}-1)-T_{1}$ resource consumption. When the algorithm resolves the fluid model at time $T_{2}$, we can derive the solution $\boldsymbol{u}^{T_{2}}$ with $u_{2}^{T_{2}}=\max\left\{\frac{1}{2}T-\nu-\frac{1}{2}(T-T_{2}+1),0\right\}\leq\max\left\{\frac{1}{2}(T_{2}-1)-\frac{1}{2}(T_{2}-1)-T_{1},0\right\}\leq T_{1}$. Therefore, the algorithm will accept at most $\frac{3}{2}T_{1}$ type-2 customers and induce at least $(1-\sqrt{c_{12}})\sqrt{T}-\frac{3}{2}T_{1}=\Omega(\sqrt{T})$ revenue loss under this event. Then, we provide a lower bound for the event probability as follows (Let $\Phi(\cdot)$ denote the c.d.f of a standard normal distribution):
$\mathbb{P}\left(\frac{1}{2}(T_{2}-1)\leq\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)+\sqrt{T_{2}-1}\;\&\;\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)-\sqrt{T}\right)$
$=$ $\mathbb{P}\left(0\leq\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq 2\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq-\frac{\sqrt{T}}{\frac{1}{2}\sqrt{T-T_{2}+1}}\right)$
$\geq$ $\mathbb{P}\left(0\leq\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq 2\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq-\frac{2}{\sqrt{1-c_{12}}}\right)$
$\geq$ $\left(0.477-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right)\cdot\left(\Phi\left(-\frac{2}{\sqrt{1-c_{12}}}\right)-\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{2}+1}}\right),$
which is greater than a constant when $T$ is above a threshold. Therefore, the regret of the algorithm is at least $\Omega(\sqrt{T})$.
Case II: $T-T_{2}+1=\Omega(\sqrt{T})$ and $T-T_{2}+1=o(T)$.
In this case, we consider the event when $\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T-T_{2}+1}$ and $\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}\leq\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)+\sqrt{T-T_{2}+1}$. Under this event, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}T$, and hence the hindsight optimal policy accepts all type-1 customers. Then, we analyze the performance of the AIR-KP policy with the resolving schedule $\{T_{1},T_{2}\}$. Similar to Case I, the algorithm rejects all customers during the interval $[1,T_{1})$. During the interval $[T_{1},T_{2})$, the algorithm accepts all type-1 customers and at most $\frac{1}{2}T_{1}$ type-2 customers. Then, the algorithm resolves the fluid model and get the solution $\boldsymbol{u}^{T_{2}}$ with $u_{1}^{T_{2}}=\min\left\{\frac{1}{2}T-\nu,\frac{1}{2}(T-T_{2}+1)\right\}\leq\frac{1}{2}(T-T_{2}+1)$. However, since $\Lambda_{1}(T_{2},T)\geq\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}$, the algorithm rejects at least $\frac{1}{2}\sqrt{T-T_{2}+1}$ type-1 customers, resulting in $\Omega(T^{1/4})$ revenue loss. Then, given a sufficiently large $T$, we provide lower bound for the event probability as follows:
$\mathbb{P}\left(\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T-T_{2}+1}\;\&\;\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}\leq\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)+\sqrt{T-T_{2}+1}\right)$
$=$ $\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-\frac{\sqrt{T-T_{2}+1}}{\frac{1}{2}\sqrt{T_{2}-1}}\right)\cdot\mathbb{P}\left(1\leq\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq 2\right)$
$\geq$ $\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-2\right)\cdot\mathbb{P}\left(1\leq\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq 2\right)$
$\geq$ $\left(0.022-\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right)\cdot\left(0.135-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{2}+1}}\right),$
which is greater than a constant when $T$ is above a threshold. Therefore, the regret of the algorithm is at least $\Omega(T^{1/4})$.
Case III: $T-T_{2}+1=o(\sqrt{T})$.
In this case, we consider the event when $\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}$. Under this event, given a sufficiently large $T$, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}+(T-T_{2}+1)\leq\frac{1}{2}(T_{2}-1)-\frac{1}{2}\sqrt{T_{2}-1}$, and hence the hindsight optimal policy accepts at least $\frac{1}{2}\sqrt{T_{2}-1}$ type-2 customers. Then, similar to the analysis in Case I, the algorithm rejects all customers during the interval $[1,T_{1})$. During the interval $[T_{1},T_{2})$, the algorithm accepts all type-1 customers and at most $\frac{1}{2}T_{1}$ type-2 customers. During the interval $[T_{2},T]$, the algorithm accepts at most $T-T_{2}+1$ type-2 customers. Thus, the algorithm accepts at most $\frac{1}{2}T_{1}+T-T_{2}+1$ type-2 customers, resulting in $\frac{1}{2}\sqrt{T_{2}-1}-\frac{1}{2}T_{1}-(T-T_{2}+1)=\Omega(\sqrt{T})-\mathcal{O}(T^{1/4})-o(T^{1/4})=\Omega(\sqrt{T})$. Then, we provide a lower bound for the event probability as follows:
$\mathbb{P}\left(\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}\right)=\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-2\right)\geq\left(0.022-\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right),$
which is greater than a constant when $T$ is greater than a threshold. Therefore, the regret will be $\Omega(\sqrt{T})$.
To summarize, given two resolving time points, the regret of the AIR-KP policy is $\Omega(T^{1/4})$. ∎
B.14 Proof of Lemma 3